OpenStudy (anonymous):

solve 4^(x + 7) = 6^(x-1)

OpenStudy (klimenkov):

\(4^{x+7}=6^{x-1}\) \(4^7\cdot4^{x}=\frac16\cdot6^{x}\) \(\left(\frac46\right)^{x}=\frac1{6\cdot4^7}\) \(x=\log_\frac46\frac1{6\cdot4^7}\)